### v2

# Minimizing Selection Bias in Inclusive Jets in Heavy-Ion Collisions with Energy Correlators — arXiv note Carlota Andres

Abstract
This note serves as a companion to a Letter Andres et al. (2024a), where we introduce a new energy correlator-based observable designed to minimize the impact of selection bias due to energy loss in inclusive jets in heavy-ion collisions. Here, we apply the method outlined in the Letter to the first-ever measurement of energy correlators in heavy-ion collisions, recently released by the CMS Collaboration Viinikainen (2024); CMS (2024).
††preprint: MIT-CTP/5764
This brief note applies the method introduced in Andres et al. (2024a) to the CMS measurement of the two-point energy correlator (E2C) in inclusive jets in heavy-ion (A-A) collisions Viinikainen (2024); CMS (2024). In Andres et al. (2024a), we introduced a new correlator-based observable, defined as
where denotes the E2C distribution Chen et al. (2020); Komiske et al. (2023) in heavy-ion jets, with being the angle between detectors, and is the so-called unbiasing function, designed to correct for leading-order energy loss effects the E2C distribution. For the derivation and explicit functional form of the unbiasing function, we refer the reader to Andres et al. (2024a). Once the E2C distribution is measured, the unbiasing function can be directly obtained. We note that in proton-proton (p-p) collisions, is unity by construction, making identical to E2C for p-p jets.
The CMS measurement uses data in lead-lead (Pb- Pb) collisions at center-of-mass energy TeV. The E2C distribution was measured for charged particles on inclusive jets with transverse momenta between GeV and GeV reconstructed using the anti- algorithm with . Once a jet is reconstructed, the winner takes all axis was found and a cone was drawn around it. All the particles measured inside that cone were correlated. A vacuum reference was analogously found using data on p-p collisions at the same center of mass energy CMS (2024). In the following, we will compare the E2C and E2C/ across different reconstructed jet transverse momenta and centrality classes. It is important to note that the error bands on the E2C/ curves we present are likely overestimated. Since the unbiasing function is derived directly from the E2C distribution, its uncertainties are expected to be highly correlated with those of the E2C itself. However, our current analysis does not account for these correlations in the error propagation. A comprehensive experimental analysis will undoubtedly refine and improve upon the results presented in this Note.
In Fig. 1, we present a comparison of the E2C (left panel) and E2C/ (right panel) for inclusive jets with reconstructed in the range GeV. Results are shown for both p-p and the 0-10 centrality class of Pb-Pb collisions. In the top left panel, we observe that the Pb-Pb E2C (green) is shifted toward smaller values compared to the p-p E2C (black). This shift results in the overall suppression of the Pb-Pb spectrum relative to the p-p one for , as illustrated in the bottom left panel. When the unbiasing function is applied (right panel), the suppression due to energy loss is significantly reduced. Moreover, the ratio between the Pb-Pb E2C/ and the p-p one (right bottom panel) becomes completely flat in the hadronization region (). This shows that, within the current uncertainties, the observed modifications in the E2C in Pb-Pb compared to p-p in the deep hadronization region (left panel, ) are fully attributable to selection bias effects. In the large angular region (), we observe that the enhancement with respect to p-p is larger for the E2C/ distribution than for the E2C, indicating that the application of the unbiasing function has effectively mitigated energy loss effects in this region. However, a quantitative description of the remaining enhancement in the E2C/ spectrum requires further investigation, as several phenomena—such as medium-induced radiation and medium response —are predicted to contribute to this behavior Andres et al. (2023a, b, 2024b); Bossi et al. (2024); Yang et al. (2024).
Figure 1: Left panel: The top plot shows the E2C for inclusive anti- jets with reconstructed in the range GeV within the 0-10 centrality class in TeV Pb-Pb collisions (green), compared to that of inclusive jets in TeV collisions (black) in the same range. The bottom plot shows the ratio of the E2C for Pb-Pb jets to that of p-p jets. Right panel: The top plot shows the E2C/ for inclusive jets with reconstructed in the range GeV within the 0-10 centrality class in TeV Pb-Pb collisions (blue), compared to the E2C of inclusive jets in TeV collisions (black) in the same range. The bottom plot shows the ratio of the E2C/ for Pb-Pb jets to the E2C of p-p jets. We note that for p-p jets.
In Fig. 2, we analyze the centrality dependence of the E2C and E2C/ spectra. The filled squares (open circles) represent the E2C (E2C/) within inclusive jets with reconstructed in the range GeV divided by the corresponding p-p baseline. Each panel corresponds to a different centrality class in Pb-Pb collisions. It is evident that as centrality decreases, the amplitude of the wide-angle enhancement in both the E2C and E2C/ ratios diminishes. This observation aligns with the general prediction that both selection bias due to energy loss and other medium effects are reduced in more peripheral collisions Andres et al. (2023a, b, 2024b); Bossi et al. (2024); Viinikainen (2024). The - centrality class is particularly noteworthy. Before applying the unbiasing function (filled dark blue squares), the Pb-Pb/p-p ratio within this class appears to display a nontrivial structure. While, after applying the unbiasing function, the E2C/ Pb-Pb/p-p ratio (empty light blue circles) becomes completely flat. This indicates that a small selection bias was the sole cause of deviation from the vacuum baseline and that the unbiasing function has accurately corrected for this effect.
Figure 2: Top left: E2C (full squares) and E2C/ (empty circles) for inclusive anti- jets with reconstructed in the range GeV within the 0-10 centrality class in TeV Pb-Pb collisions (squares) divided by the E2C of inclusive jets in TeV p-p collisions in the same range.
Top right: As in the top left, but for jets in 10-30 Pb-Pb collisions. Bottom left: As in the top left, but for jets in 30-50 Pb-Pb collisions. Bottom right: As in the top left, but for jets in 50-90 Pb-Pb collisions.
Fig. 3 focuses on the -dependence of the results. The filled squares (empty circles) represent the E2C (E2C/) distribution within inclusive jets in the 0-10 centrality class in Pb-Pb collisions divided by the E2C distribution in inclusive p-p jets. Each panel corresponds to a different range of reconstructed jet transverse momenta. It can be clearly observed that the wide-angle enhancement decreases in amplitude as the jet increases. Additionally, the wide-angle structure shifts to slightly smaller values as the jet increases. These qualitative behaviors were predicted as consequences of medium-induced radiation prior to the unveiling of the CMS measurement Andres et al. (2023a, b, 2024b). However, a comprehensive analysis, including the extraction of scaling laws governing this enhancement, is required to disentangle medium-induced radiation from other effects, such as medium response Yang et al. (2024); Bossi et al. (2024), and to identify possible signatures of color coherence dynamics Andres et al. (2023a, b).
Figure 3: Top left: E2C (full squares) and E2C/ (empty circles) for inclusive anti- jets with reconstructed in the range GeV within the 0-10 centrality class in TeV Pb-Pb collisions (squares) divided by the E2C of inclusive jets in TeV p-p collisions in the same range.
Top right: As in the top left, but for jets with in the range 140–160 GeV. Bottom left: As in the top left, but for jets with in the range 160–180 GeV. Bottom right: As in the top left, but for jets with in the range 180–200 GeV.
Finally, we present in Fig. 4 the evaluation of the following approximate formula, derived in Andres et al. (2024a),
using the CMS p-p data on the two-point energy correlator within p-p jets CMS (2024) and for . We refer the interested reader to Andres et al. (2024a) for further details on this equation and the role it plays in the derivation of the unbiasing function . The derivative on the right hand side of eq. (2) was evaluated analytically, while that on the left hand side was computed numerically. We did not compute error bars for the numerical derivatives, as the bin-to-bin correlated uncertainties of the CMS measurement are not yet available. The results shown in Fig. 4 show good agreement between both sides of eq. (2) over the complete angular range studied, even around the hadronization transition at .
Figure 4: Direct evaluation of Eq. (2) for using the experimental data on the E2C within inclusive anti- jets in TeV p-p collisions CMS (2024). The left hand side of Eq. (2) is represented by solid curves, while the right-hand side is shown by empty markers. Different colors correspond to results for different reconstructed jet ranges. For clarity, offsets of 0.3, 0.6, and 0.9 have been applied to the results for jets with ranges of GeV, GeV, and GeV ranges, respectively. Note that, as explained in the main text, error bands are not displayed on the curves representing the l.h.s. of eq. (2).
Conclusions—This companion note to the Letter Andres et al. (2024a) presents results on the correlator-based observable introduced in Andres et al. (2024a), E2C/, using the CMS measurement of the two-point energy (E2C) correlator within inclusive jets in Pb-Pb collisions at the LHC Viinikainen (2024); CMS (2024). The E2C/ observable shows significantly reduced sensitivity to energy loss effects compared to the corresponding E2C for jets with transverse momenta between 120 GeV and 200 GeV, across different centrality classes in Pb-Pb collisions. Our current analysis does not account for correlations in the error propagation, and thus we encourage the experiments to perform this measurement.
Acknowledgements.— We are especially grateful to Raghav Kunnawalkam Elayavalli and Jussi Viinikainen for helping throughout this project and for granting us the privilege of being the sole authors of this note. We thank Hannah Bossi, Fabio Dominguez, Kyle Lee, Yen-Jie Lee, Cyrille Marquet, Ian Moult, Krishna Rajagopal, and Carlos A. Salgado for useful discussions. The work of CA was partially supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under grant Contract Number DE-SC0011090 and by OE Portugal, Fundação para a Ciência e a Tecnologia (FCT), I.P., project 2024.06117.CERN. CA also acknowledges financial support by the FCT under contract 2023.07883.CEECIND. The authors would like to express special thanks to the Mainz Institute for Theoretical Physics (MITP) of the Cluster of Excellence PRISMA+ (Project ID 390831469), for its hospitality and support.